Suppose an atom of Osmium at rest emits an X-ray photon of energy 6.8 keV. Calculate the
“recoil” momentum and kinetic energy of the atom.
1
Expert's answer
2020-06-15T09:48:53-0400

Since, system has initial momentum is zero and no external interaction is going on with the system,Momentum of the system is conserved, hence, by conservation law of momentum

"p_{osmium}+p_{\gamma}=0"

We have given that, emitted photon has energy is

"8.6keV=8.6\times 10^3eV"

Thus,

"E_{\gamma}=8.6\times 10^3eV=\frac{hc}{\lambda}=\frac{h}{\lambda}c=p_{\gamma}c\\
\implies p_{\gamma}=\frac{E_{\gamma}}{c}=8.6keV/c"

Thus,

"p_{osmium}=-8.6keV\/c"

Kinetic energy is

"K=\frac{p^2}{2m}"

Hence,

"K_{osmium}=2.088\times 10^{-4}eV"